Suppose a countrys debt is $630 billion before Year 1. Using the information below, fill in the blanks for its budget balance and debt for the following years: 1. Expenditures Debt $202 Revenues Year 1 $203 Year 2 $215 Year 3 $227 Year 4 Year 5 $258 $626 $221 $230 $243 $619
Answer #1

change in Debt = Expenditures - Revenues

Year 1 = Debt = 630 - 1 = 629

Year 2 = Expenditures = 215 + 3 = 218

Year 3 = Debt = 626 - 6 = 620

Year 4 = Revenues = 230 + 1 = 231

Year 5 = Debt 619 - 15 = 604